OpenStudy (ayanaortega):

What are the solutions to the equation? Use the quadratic formula to solve the equation. 2y^2-6y-8=0

MsBrains (ms-brains):

Let's solve your equation step-by-step. 2y2−6y−8=0 Step 1: Factor left side of equation. 2(y+1)(y−4)=0 Step 2: Set factors equal to 0. y+1=0 or y−4=0 y=−1 or y=4 Answer: y=−1 or y=4 Sources: Mathpapa.com

OpenStudy (ayanaortega):

4,-1

MsBrains (ms-brains):

yes